Determine the dividends received deduction in each of the following situations. Assume that none of the corporate shareholders owns 20% or more of the corporations paying the dividends.

Red

Purple

Orange

Income from operations

$450,000

$550,000

$550,000

Expenses from operations

350,000

600,000

560,000

Qualifying dividends

100,000

100,000

100,000
Answer #1

a. The dividends received deduction for Red Corporation is $50000

b. The dividends received deduction for Purple Corporation is $50000

c. The dividends received deduction for Orange Corporation is $45000

Red

purple

Orange

Taxable Income (before DRD)

200000

50000

90000

Step 1: (dividend * deduction %)

50000

50000

50000

Step 2: (tax income* deduction %)

100000

25000

45000

Step 3: (deduct lesser of step 1 or step 2, UNLESS NOL rule applies)

200000-50000 = Positive, the value of step 2 should be deducted, that is, $50000

50000-50000= 0

NOL rule applies.

Therefore, entire $50000 should be deducted

90000-50000= positive, the value of step 2 should be deducted, that is, $45000